# Force-free wave interaction in magnetar magnetospheres: Computational modeling in axisymmetry

### v1

@@ Abstract
Crustal quakes of highly magnetized neutron stars can disrupt their magnetospheres, triggering energetic phenomena like X-ray and fast radio bursts (FRBs). Understanding plasma wave dynamics in these extreme environments is vital for predicting energy transport across scales to the radiation length. This study models relativistic plasma wave interaction in magnetar magnetospheres with force-free electrodynamics simulations. For propagation along curved magnetic field lines, we observe the continuous conversion of Alfvén waves (AWs) to fast magnetosonic (FMS) waves. The conversion efficiency can be up to three times higher when counter-propagating AWs interact in the equatorial region. Alfvén waves generate FMS waves of twice their frequency during their first crossing of the magnetosphere. After the initial transient burst of FMS waves, AWs convert to FMS waves periodically, generating variations on timescales of the magnetospheric AW crossing time. This decaying FMS wave tail carries a significant portion (half) of the total energy emitted. Plastic damping of ‘bouncing’ AWs by the magnetar crust has minimal impact on the FMS efficiency. We discuss the implications of the identified wave phenomena for magnetar observations. Outgoing FMS waves can develop electric zones, potential sources of coherent radiation. Long wavelength FMS waves could generate FRBs through reconnection beyond the light cylinder.
@@ 1 Introduction
Magnetars are neutron stars with ultra-strong magnetic fields of $10^{14}-10^{16}\,{\rm G}$ (see reviews by Kaspi & Beloborodov, 2017; Esposito et al., 2020). They exhibit rich radiative activities. Transient X-ray events include powerful giant flares with luminosities up to $10^{47}\,{\rm erg/s}$ on millisecond timescales, and lower energy outbursts of $10^{35}\,{\rm erg/s}$ with month-to-year long decays. Quasi-periodic oscillations (QPOs) with frequencies ranging from approximately $150\,{\rm Hz}$ to $500\,{\rm Hz}$ are commonly observed during giant flares (Barat et al., 1983; Strohmayer & Watts, 2005; Israel et al., 2005; Castro-Tirado et al., 2021). At least in some cases, these QPOs have been attributed to global torsional vibrations (e.g. Strohmayer & Watts, 2006; Levin, 2007; Gabler et al., 2011). The energy source of these events is likely the magnetic field. Sudden global magnetospheric rearrangement could power fast giant flares (Thompson & Duncan, 1995, 2001; Parfrey et al., 2013; Carrasco et al., 2019b; Mahlmann et al., 2022), while slower dissipation of magnetospheric twist was proposed to explain persistent magnetar emissions (Beloborodov, 2013). In addition, the galactic magnetar SGR 1935+2154 is a source of fast radio bursts (FRBs, CHIME/FRB Collaboration, 2020). FRB polarization and timing information offers insights into the highly magnetized environments of compact astrophysical objects. Plasma instabilities in the magnetized environment of magnetars offer a suitable energy reservoir to power the coherent generation of FRBs (reviews, e.g., in Lyubarsky, 2021; Petroff et al., 2022; Zhang, 2023). How magnetic energy is dissipated in the magnetosphere to power X-ray bursts and FRBs is the key theoretical question that remains to be answered.
@@
One possible mechanism of energy dissipation in the highly relativistic magnetar magnetosphere is nonlinear wave interactions (e.g., Lyubarsky, 2018; TenBarge et al., 2021; Ripperda et al., 2021; Golbraikh & Lyubarsky, 2023). Starquakes (Thompson & Duncan, 2001) and crustal failures in the stellar interior (Li et al., 2016; Thompson et al., 2017) can perturb the magnetosphere by displacing footpoints of magnetic field lines and excite waves that carry magnetic energy into the extended magnetosphere. Waves in highly magnetized plasmas can be modeled well by force-free electrodynamics (FFE), assuming a vanishing Lorentz force with plasma particles streaming along magnetic field lines (Uchida, 1997; Gruzinov, 1999; Komissarov, 2002). FFE supports two wave modes as solutions to the full nonlinear FFE equations: Alfvén waves (AWs) and fast magnetosonic (FMS) waves. Alfvén waves transport energy along the direction of the background field while FMS waves propagate similarly to waves in vacuum electrodynamics – across magnetic field lines. Thompson & Duncan (2001) and Thompson (2022) proposed that the nonlinear interaction between Alfvén waves can drive turbulent cascades to small scales that heat the magnetospheric pair plasma to form a radiatively efficient ‘fireball’ that explains magnetar X-ray bursts.
@@
The theory of wave interactions and their resulting spectra is long debated. Sridhar & Goldreich (1994) studied weak Alfvén turbulence in non-relativistic incompressible fluids and found leading order four-wave interactions resulting in an anisotropic spectrum. Three-wave interactions are kinetically prohibited (only possible if one of the involved Alfvén waves is non-propagating, see Ng & Bhattacharjee, 1996; Goldreich & Sridhar, 1997). FMS waves can be excited to enable three-wave interactions in compressible fluids and play an important role in the energy transfer (Thompson & Blaes, 1998; Takamoto & Lazarian, 2016, 2017). Similar conclusions are also found in relativistic turbulence (Thompson & Blaes, 1998; Cho, 2005; Zrake & MacFadyen, 2012; Zrake & East, 2016; Takamoto & Lazarian, 2016, 2017; Li et al., 2019; TenBarge et al., 2021; Ripperda et al., 2021). Numerical models (Li et al., 2019; Mahlmann et al., 2020a; Ripperda et al., 2021; Nättilä & Beloborodov, 2022) of relativistic Alfvén turbulence suggest that the turbulent dissipation may be slow compared to competing processes like Alfvén wave absorption by the stellar surface (Li & Beloborodov, 2015). However, these works study Alfvén wave interactions in local settings with uniform background magnetic fields. AW dynamics in dipole magnetospheres tied to a stellar surface require more complex models.
@@
Recently, nonlinear wave dynamics have been playing a prominent role in various FRB models (Lyubarsky, 2020; Yuan et al., 2020; Lu et al., 2020). Alfvén waves increase in amplitudes compared to the decaying magnetospheric fields as they propagate to the outer magnetosphere. When they become non-linear, they open up the magnetosphere to form plasmoids, ignite magnetic reconnection, and drive relativistic blast waves (Yuan et al., 2020) that could eventually excite radio emission by the synchrotron maser instability (e.g., Ghisellini, 2016; Plotnikov & Sironi, 2019; Metzger et al., 2019). Alfvén waves propagating along curved magnetic fields are a source of FMS waves due to non-linear coupling with the background magnetic field (Yuan et al., 2021; Chen et al., 2024; Bernardi et al., 2024). Long wavelength FMS pulses are relevant for FRB models because they can perturb the current sheet beyond the Y-point or in the magnetar wind to generate radio emission (Lyubarsky, 2020; Mahlmann et al., 2023; Wang et al., 2023). In this context, it is crucial to disentangle all the different processes involved in the Alfvén wave dynamics: wave interactions, cascading of energy from large scales to small scales, mode conversion and dissipation.
@@
In this work, we study essential aspects of Alfvén wave interactions in axisymmetric dipolar magnetospheres to derive limits for FMS wave generation enhanced by non-linear interactions. This paper is organized as follows. Section 2 reviews the basic theory of FFE magnetospheric modeling and the possible wave interactions (Section 2.1). We revisit the mode conversion at curved field lines (Section 2.2) and through non-linear interactions (Section 2.3). Section 3 describes our basic numerical setup (Section 3.1) and analyzes various FFE simulations: Section 3.2 evaluates the conversion of long wave packets of AWs to FMS waves in the inner magnetosphere. Section 3.3 shows how AW collisions enhance FMS wave generation, and Section 3.4 analyzes the long-term dynamics of counter-propagating AWs. We discuss the implications of our results in Section 4 and summarize our conclusions in Section 5.
@@ 2 Theoretical background
Figure 1: Approximate scaling of the coupling strength between Alfvén waves propagating along curved field lines and the FMS wave electric field (Equation 16; $b_{\rm eq}=1$). Black dots indicate colatitudes at which field lines are anchored to the star (red line: equator). AWs injected in the northern hemisphere propagate along dipolar field lines through the magnetosphere (varying their $\theta$ position).
Deformations of the magnetar crust can generate Alfvén waves propagating along magnetic field lines with wavelengths comparable to a significant fraction of the stellar radius $R_{\ast}$ (e.g., Thompson et al., 2017). On such scales, magnetic field dynamics of the highly relativistic magnetar magnetospheres can be described by force-free electrodynamics (FFE, see Blandford, 2002; Gruzinov, 1999; Spitkovsky, 2006), the high-magnetization limit of relativistic ideal magnetohydrodynamics (MHD). In this limit, electric ($\mathbf{E}$) and magnetic ($\mathbf{B}$) fields follow from Maxwell’s equations:
@@
Equations (1) to (5) allow for two wave modes: Fast magnetosonic waves with $\omega=\pm ck$, and Alfvén waves with $\omega=\pm ck_{\parallel}$. Here, $\omega$ denotes the wave frequency, and $\mathbf{k}$ is the wave vector with norm $k$. Vectors can be decomposed into projections parallel ($\parallel$) and perpendicular ($\perp$) to the magnetic field. Detailed reviews of FFE waves were presented, for example, by Punsly (2003) and Li et al. (2019). The following sections recap the properties as well as linear and non-linear wave dynamics of force-free dipolar magnetar magnetospheres.
@@ 2.1 Waves in the dipole magnetosphere
Here, $\mu=B_{*}R^{3}_{*}/2$ is the magnetic moment of the star, and $B_{*}$ is the surface magnetic field strength. Displacing magnetic field lines in the toroidal direction with an amplitude $\epsilon_{0}$ at the stellar surface induces AW magnetic fields of relative strength $b=\delta B/B_{\rm dip}$. At the stellar surface, we find [^1] In axisymmetric dipole magnetospheres, Alfvén waves have a purely toroidal magnetic field $\mathbf{B}_{\rm A}=\delta B\mathbf{e}_{\phi}$, and the wave electric field has purely poloidal components. The inductive (ideal) electric field amplitude is $\mathbf{E}_{\rm A}=-(\boldsymbol{\Omega}/c\times\mathbf{R})\times\mathbf{B}_{\rm dip}$ for purely toroidal shear velocities $v^{\phi}=\Omega R_{\ast}/c$. Inserting the dipolar field components from Equation (6) allows us to compute the wave amplitude $\delta E^{2}=\mathbf{E}_{\rm A}\cdot\mathbf{E}_{\rm A}$. We approximate the corresponding magnetic field amplitude by assuming $\delta B_{A}\approx\delta E_{A}$.:
@@
Thus, the relative strength follows $b\propto B_{\rm dip}^{-1/2}\propto r^{3/2}$. We define the critical non-linearity radius $\tilde{r}$ as the radius where $b=1$, such that
@@
In any of the previous expressions, we label field lines anchored to the star by their radius of equatorial crossing, $R_{\rm eq}=R_{*}/\sin^{2}\theta_{s}$. The relative strength of the magnetic perturbation at the equator is $b_{\rm eq}=b_{0}\sin^{-3}\theta_{s}$. FMS waves propagate like vacuum electrodynamic modes in FFE. As they can cross magnetic field lines, their dynamics on a background magnetic field often resemble concentric waves like those spreading on the surface of calm water. In the axisymmetric dipole magnetic field, the polarization of FMS waves requires toroidal wave electric fields $\delta\mathbf{E}_{F}=E\mathbf{e}_{\phi}$ (see detailed reviews by, e.g., Li et al., 2019). As FMS waves have magnetic fields parallel or antiparallel to the background magnetic field, they can develop zones where the electric field nearly (or, even completely) dominates. In these ‘electric zones’ the wave electric field becomes comparable to the total magnetic field (Beloborodov, 2023; Chen et al., 2022b; Levinson, 2022). The location where electric zones develop on the rapidly decaying dipole field ($B\propto r^{-3}$) depends on the colatitude, the injection location, and the injection amplitude. For an injection of an FMS wave at the stellar surface, we estimate the non-linearity radius:
@@
This estimate follows the derivation by Beloborodov (2023, Equation 110). Once electric zones develop, they can drive plasma heating as well as shock formation (Beloborodov, 2023; Chen et al., 2022b). We discuss the observational implications for the scenario of magnetospheric AW collisions in Section 4.

### v2

@@ Abstract
Crustal quakes of highly magnetized neutron stars can disrupt their magnetospheres, triggering energetic phenomena like X-ray and fast radio bursts (FRBs). Understanding plasma wave dynamics in these extreme environments is vital for predicting energy transport across scales to the radiation length. This study models relativistic plasma wave interaction in magnetar magnetospheres with force-free electrodynamics simulations. For propagation along curved magnetic field lines, we observe the continuous conversion of Alfvén waves to fast magnetosonic (FMS) waves. The conversion efficiency can be up to three times higher when counter-propagating Alfvén waves interact in the equatorial region. Alfvén waves generate FMS waves of twice their frequency during their first crossing of the magnetosphere. After the initial transient burst of FMS waves, Alfvén waves convert to FMS waves periodically, generating variations on timescales of the magnetospheric Alfvén wave crossing time. This decaying FMS wave tail carries a significant portion (half) of the total energy emitted. Plastic damping of ‘bouncing’ Alfvén waves by the magnetar crust has minimal impact on the FMS efficiency. We discuss the implications of the identified wave phenomena for magnetar observations. Outgoing FMS waves can develop electric zones, potential sources of coherent radiation. Long-wavelength FMS waves could generate FRBs through reconnection beyond the light cylinder.
@@ 1 Introduction
Magnetars are neutron stars with ultrastrong magnetic fields of $10^{14}-10^{16}\,{\rm G}$ (see reviews by Kaspi & Beloborodov, 2017; Esposito et al., 2020). They exhibit rich radiative activities. Transient X-ray events include powerful giant flares with luminosities up to $10^{47}\,{\rm erg/s}$ on millisecond timescales, and lower energy outbursts of $10^{35}\,{\rm erg/s}$ with month-to-year long decays. Quasiperiodic oscillations (QPOs) with frequencies ranging from approximately $150\,{\rm Hz}$ to $500\,{\rm Hz}$ are commonly observed during giant flares (Barat et al., 1983; Strohmayer & Watts, 2005; Israel et al., 2005; Castro-Tirado et al., 2021). At least in some cases, these QPOs have been attributed to global torsional vibrations (e.g. Strohmayer & Watts, 2006; Levin, 2007; Gabler et al., 2011). The energy source of these events is likely the magnetic field. Sudden global magnetospheric rearrangement could power fast giant flares (Thompson & Duncan, 1995, 2001; Parfrey et al., 2013; Carrasco et al., 2019b; Mahlmann et al., 2022), while slower dissipation of magnetospheric twist was proposed to explain persistent magnetar emissions (Beloborodov, 2013). In addition, the galactic magnetar SGR 1935+2154 is a source of fast radio bursts (FRBs; CHIME/FRB Collaboration, 2020). FRB polarization and timing information offers insights into the highly magnetized environments of compact astrophysical objects. Plasma instabilities in the magnetized environment of magnetars offer a suitable energy reservoir to power the coherent generation of FRBs (see reviews, e.g., in Lyubarsky, 2021; Petroff et al., 2022; Zhang, 2023). How magnetic energy is dissipated in the magnetosphere to power X-ray bursts and FRBs is the key theoretical question that remains to be answered.
@@
One possible mechanism of energy dissipation in the highly relativistic magnetar magnetosphere is nonlinear wave interactions (e.g., Lyubarsky, 2018; TenBarge et al., 2021; Ripperda et al., 2021; Golbraikh & Lyubarsky, 2023). Starquakes (Blaes et al., 1989; Thompson & Duncan, 2001) and crustal failures in the stellar interior (Li et al., 2016; Thompson et al., 2017) can perturb the magnetosphere by displacing footpoints of magnetic field lines and excite waves that carry magnetic energy into the extended magnetosphere. Waves in highly magnetized plasmas can be modeled well by force-free electrodynamics (FFE), assuming a vanishing Lorentz force with plasma particles streaming along magnetic field lines (Uchida, 1997; Gruzinov, 1999; Komissarov, 2002). FFE supports two wave modes as solutions to the full nonlinear FFE equations: Alfvén waves and fast magnetosonic (FMS) waves. Alfvén waves transport energy along the direction of the background field while FMS waves propagate similarly to waves in vacuum electrodynamics – across magnetic field lines. Thompson & Duncan (2001) and Thompson (2022) proposed that the nonlinear interaction between Alfvén waves can drive turbulent cascades to small scales that heat the magnetospheric pair plasma to form a radiatively efficient ‘fireball’ that explains magnetar X-ray bursts.
@@
The theory of wave interactions and their resulting spectra is long debated. Sridhar & Goldreich (1994) studied weak Alfvén turbulence in nonrelativistic incompressible fluids and found leading order four-wave interactions resulting in an anisotropic spectrum. Three-wave interactions are kinetically prohibited (only possible if one of the involved Alfvén waves is nonpropagating, see Ng & Bhattacharjee, 1996; Goldreich & Sridhar, 1997). FMS waves can be excited to enable three-wave interactions in compressible fluids and play an important role in the energy transfer (Thompson & Blaes, 1998; Takamoto & Lazarian, 2016, 2017). Similar conclusions are also found in relativistic turbulence (Thompson & Blaes, 1998; Cho, 2005; Zrake & MacFadyen, 2012; Zrake & East, 2016; Takamoto & Lazarian, 2016, 2017; Li et al., 2019; TenBarge et al., 2021; Ripperda et al., 2021). Numerical models (Li et al., 2019; Mahlmann et al., 2020a; Ripperda et al., 2021; Nättilä & Beloborodov, 2022) of relativistic Alfvén turbulence suggest that the turbulent dissipation may be slow compared to competing processes like Alfvén wave absorption by the stellar surface (Li & Beloborodov, 2015). However, these works study Alfvén wave interactions in local settings with uniform background magnetic fields. Alfvén wave dynamics in dipole magnetospheres tied to a stellar surface require more complex models.
@@
Recently, nonlinear wave dynamics have been playing a prominent role in various FRB models (Lyubarsky, 2020; Yuan et al., 2020; Lu et al., 2020). Alfvén waves increase in amplitudes compared to the decaying magnetospheric fields as they propagate to the outer magnetosphere. When they become nonlinear, they open up the magnetosphere to form plasmoids, ignite magnetic reconnection, and drive relativistic blast waves (Yuan et al., 2020) that could eventually excite radio emission by the synchrotron maser instability (e.g., Ghisellini, 2016; Plotnikov & Sironi, 2019; Metzger et al., 2019). Alfvén waves propagating along curved magnetic fields are a source of FMS waves due to nonlinear coupling with the background magnetic field (Yuan et al., 2021; Chen et al., 2024; Bernardi et al., 2024). Long-wavelength FMS pulses are relevant for FRB models because they can perturb the current sheet beyond the Y-point or in the magnetar wind to generate radio emission (Lyubarsky, 2020; Mahlmann et al., 2023; Wang et al., 2023). In this context, it is crucial to disentangle all the different processes involved in the Alfvén wave dynamics: wave interactions, cascading of energy from large scales to small scales, mode conversion, and dissipation.
@@
In this work, we study essential aspects of Alfvén wave interactions in axisymmetric dipolar magnetospheres to derive limits for FMS wave generation enhanced by nonlinear interactions. This paper is organized as follows. Section 2 reviews the basic theory of FFE magnetospheric modeling and the possible wave interactions (Section 2.1). We revisit the mode conversion at curved field lines (Section 2.2) and through nonlinear interactions (Section 2.3). Section 3 describes our basic numerical setup (Section 3.1) and analyzes various FFE simulations: Section 3.2 evaluates the conversion of long wave packets of Alfvén waves to FMS waves in the inner magnetosphere. Section 3.3 shows how Alfvén wave collisions enhance FMS wave generation, and Section 3.4 analyzes the long-term dynamics of counter-propagating Alfvén waves. We discuss the implications of our results in Section 4 and summarize our conclusions in Section 5.
@@ 2 Theoretical background
Figure 1: Approximate scaling of the coupling strength between Alfvén waves propagating along curved field lines and the FMS wave electric field (Equation 16; $b_{\rm eq}=1$). Black dots indicate colatitudes at which field lines are anchored to the star (red line: equator). Alfvén waves injected in the northern hemisphere propagate along dipolar field lines through the magnetosphere (varying their $\theta$ position).
Deformations of the magnetar crust can generate Alfvén waves propagating along magnetic field lines with wavelengths comparable to a significant fraction of the stellar radius $R_{\ast}$ (e.g., Thompson et al., 2017). On such scales, magnetic field dynamics of the highly relativistic magnetar magnetospheres can be described by FFE (see Blandford, 2002; Gruzinov, 1999; Spitkovsky, 2006), the high-magnetization limit of relativistic ideal magnetohydrodynamics (MHD). In this limit, electric ($\mathbf{E}$) and magnetic ($\mathbf{B}$) fields follow from Maxwell’s equations:
@@
Equations (1) to (5) allow for two wave modes: FMS waves with $\omega=\pm ck$, and Alfvén waves with $\omega=\pm ck_{\parallel}$. Here, $\omega$ denotes the wave frequency, and $\mathbf{k}$ is the wavevector with norm $k$. Vectors can be decomposed into projections parallel ($\parallel$) and perpendicular ($\perp$) to the magnetic field. Detailed reviews of FFE waves were presented, for example, by Punsly (2003) and Li et al. (2019). The following sections recap the properties as well as linear and nonlinear wave dynamics of force-free dipolar magnetar magnetospheres.
@@ 2.1 Waves in the dipole magnetosphere
Here, $\mu=B_{*}R^{3}_{*}/2$ is the magnetic moment of the star, and $B_{*}$ is the surface magnetic field strength. Deformations or faults in a magnetar crust can generate both Alfvén waves and FMS waves (see Blaes et al., 1989; Thompson & Duncan, 2001; Li et al., 2016; Thompson et al., 2017). The high incompressibility of the neutron star matter and crust implies that coupling seismic waves to Alfvén waves is more likely than launching (compressive) FMS waves. Any FMS wave launched from the stellar surface moves across magnetic field lines and carries energy away from the star. This work investigates how much of the energy of Alfvén waves, produced at the footpoints of magnetic field lines and propagating along magnetic flux tubes in the inner magnetosphere, can escape the closed magnetosphere as secondary FMS waves. Displacing magnetic field lines in the toroidal direction with an amplitude $\epsilon_{0}$ at the stellar surface induces Alfvén wave magnetic fields of relative strength $b=\delta B/B_{\rm dip}$. At the stellar surface, we find [^1] In axisymmetric dipole magnetospheres, Alfvén waves have a purely toroidal magnetic field $\mathbf{B}_{\rm A}=\delta B\mathbf{e}_{\phi}$, and the wave electric field has purely poloidal components. The inductive (ideal) electric field amplitude is $\mathbf{E}_{\rm A}=-(\boldsymbol{\Omega}/c\times\mathbf{R})\times\mathbf{B}_{\rm dip}$ for purely toroidal shear velocities $v^{\phi}=\Omega R_{\ast}/c$. Inserting the dipolar field components from Equation (6) allows us to compute the wave amplitude $\delta E^{2}=\mathbf{E}_{\rm A}\cdot\mathbf{E}_{\rm A}$. We approximate the corresponding magnetic field amplitude by assuming $\delta B_{A}\approx\delta E_{A}$.:
@@
Thus, the relative strength follows $b\propto B_{\rm dip}^{-1/2}\propto r^{3/2}$. We define the critical nonlinearity radius $\tilde{r}$ as the radius where $b=1$, such that
@@
In any of the previous expressions, we label field lines anchored to the star by their radius of equatorial crossing, $R_{\rm eq}=R_{*}/\sin^{2}\theta_{s}$. The relative strength of the magnetic perturbation at the equator is $b_{\rm eq}=b_{0}\sin^{-3}\theta_{s}$. FMS waves propagate like vacuum electrodynamic modes in FFE. As they can cross magnetic field lines, their dynamics on a background magnetic field often resemble concentric waves like those spreading on the surface of calm water. In the axisymmetric dipole magnetic field, the polarization of FMS waves requires toroidal wave electric fields $\delta\mathbf{E}_{F}=E\mathbf{e}_{\phi}$ (see detailed reviews by, e.g., Li et al., 2019). As FMS waves have magnetic fields parallel or antiparallel to the background magnetic field, they can develop zones where the electric field nearly (or, even completely) dominates. In these ‘electric zones’ the wave electric field becomes comparable to the total magnetic field (Beloborodov, 2023; Chen et al., 2022b; Levinson, 2022). The location where electric zones develop on the rapidly decaying dipole field ($B\propto r^{-3}$) depends on the colatitude, the injection location, and the injection amplitude. For an injection of an FMS wave at the stellar surface, we estimate the nonlinearity radius:
@@
This estimate follows the derivation by Beloborodov (2023, Equation 110). Once electric zones develop, they can drive plasma heating as well as shock formation (Beloborodov, 2023; Chen et al., 2022b). We discuss the observational implications for the scenario of magnetospheric Alfvén wave collisions in Section 4.